OpenStudy (anonymous):

Find the constant of variation for the relation and use it to write an equation for the statement. y is a joint variation of x and z and varies inversely with w. When x = 3, z = 4, and w = 6, y is equal to 8.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

y=(k)xz and y=k/w (where k is constant) if I am understanding correctly.

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm 8=k(3)(4) }\) \(\large\color{blue}{ \rm 8=k/6 }\) you get different `k` s -:(

OpenStudy (anonymous):

thanks I think I get it now

OpenStudy (solomonzelman):

I think something is missing in the question.... the `k` s are supposed to be the same. Otherwise, why are these different variations, (joint var. and inv. variation) in a same question ? ... idk